# Maximal non-compactness of embeddings between sequence spaces

### v1

@@ 3. Lorentz sequence spaces
Let us assume that $\ell^{p,q}\not\subset c_{0}$ and take $a\in\ell^{p,q}\setminus c_{0}$. Then there exists $\varepsilon>0$, such that for all $n_{0}\in\mathbb{N}$ there exists $n\geq n_{0},\ n\in\mathbb{N}\colon|a_{n}|\geq\varepsilon$, which implies that $a_{n}^{*}\ \geq\ \varepsilon$ for $\forall n\in\mathbb{N}$. First, let us look at the case when $q<\infty$ and $p\in(0,\infty]$. Then:
@@
The statements of the following two lemmas occur in the article [10, p. 77-78]. Since, however, the verification of the statements is only sketched there, we shall include detailed proofs.
@@
Let $p\in(0,\infty]$, $q\in(0,\infty)$ and $a=\{a_{n}\}_{n=1}^{\infty}\in\ell_{p,q}$, then for all $n\in\mathbb{N}$ it holds that
@@
    Here, the second inequality follows from the fact that $(a_{i}^{*})$ is the decreasing rearrangement. The third inequality results from the Lagrange mean value theorem. Indeed, for all $i\in\mathbb{N}$, there exists $\xi\in(i-1,i)$, that satisfies for the function $f(t)=\frac{p}{q}t^{\frac{q}{p}}$, where $t\geq 0$ following equality: f^’(ξ) = ξ^qp - 1 = pq[i^qp-(i-1)^qp]. Also $i^{\frac{q}{p}-1}\geq\xi^{\frac{q}{p}-1}$, because $\frac{q}{p}\geq 1$ and $i>\xi>1$. From the estimates above we get the third inequality and 1. is proven.
@@
    Here the second inequality follows from the fact that for all $i\leq n$ it holds that $\frac{q}{p}\in[0,1)$. Therefore $\left(\frac{q}{p}-1\right)\in[-1,0)$, so we get $n^{\left(\frac{q}{p}-1\right)}\leq i^{\left(\frac{q}{p}-1\right)}$, and 2. is proven.
@@
$\ell_{p,q_{1}}\hookrightarrow\ell_{p,q_{2}},$
and for all $a\in\ell_{p,q_{1}}$ it holds that
@@
    $\|a\|_{p,q_{2}}\leq\left(\frac{q_{1}}{p}\right)^{\frac{1}{q_{1}}-\frac{1}{q_{2}}}\|a\|_{p,q_{1}}$; if $p<q_{1}$,
@@
    1. (a)
      Let $q_{2}<\infty$. From the first part of Lemma 3.2 we get:
      $\|a\|_{p,q_{2}}^{q_{2}}$ $=\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{2}}{p}-1}(a_{n}^{*})^{q_{2}-q_{1}}(a_{n}^{*})^{q_{1}}$
      $\leq\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{2}}{p}-1}\left(\frac{q_{1}}{p}\right)^{\frac{q_{2}-q_{1}}{q_{1}}}n^{-\frac{q_{2}-q_{1}}{p}}\|a\|_{p,q_{1}}^{q_{2}-q_{1}}(a_{n}^{*})^{q_{1}}$
      $=\left(\frac{q_{1}}{p}\right)^{\frac{q_{2}}{q_{1}}-1}\|a\|_{p,q_{1}}^{q_{2}-q_{1}}\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{1}}{p}-1}(a_{n}^{*})^{q_{1}}$
      $=\left(\frac{q_{1}}{p}\right)^{\frac{q_{2}}{q_{1}}-1}\|a\|_{p,q_{1}}^{q_{2}-q_{1}}\|a\|_{p,q_{1}}^{q_{1}}$
      $=\left(\frac{q_{1}}{p}\right)^{\frac{q_{2}}{q_{1}}-1}\|a\|_{p,q_{1}}^{q_{2}}.$
    2. (b)
      Now, let $q_{2}=\infty$:
      $\|a\|_{p,q_{2}}$ $=\displaystyle\sup_{n\in\mathbb{N}}n^{\frac{1}{p}}a_{n}^{*}$
      $\leq\displaystyle\sup_{n\in\mathbb{N}}n^{\frac{1}{p}}\left(\frac{q_{1}}{p}\right)^{\frac{1}{q_{1}}}n^{-\frac{1}{p}}\|a\|_{p,q_{1}}$
      $=\left(\frac{q_{1}}{p}\right)^{\frac{1}{q_{1}}}\|a\|_{p,q_{1}}.$
      Here the first inequality again follows from Lemma 3.2.
    So 1. is proven.
@@
    Now let $p\geq q_{1}$. Then, the following inequality follows from the second part of Lemma 3.2:
    $\|a\|_{p,q_{2}}^{q_{2}}$ $=\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{2}}{p}-1}(a_{n}^{*})^{q_{2}-q_{1}}(a_{n}^{*})^{q_{1}}$
    $\leq\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{2}}{p}-1}n^{-\frac{q_{2}-q_{1}}{p}}\|a\|_{p,q_{1}}^{q_{2}-q_{1}}(a_{n}^{*})^{q_{1}}$
    $=\|a\|_{p,q_{1}}^{q_{2}-q_{1}}\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{1}}{p}-1}(a_{n}^{*})^{q_{1}}$
    $=\|a\|_{p,q_{1}}^{q_{2}-q_{1}}\|a\|_{p,q_{1}}^{q_{1}}$
    $=\|a\|_{p,q_{1}}^{q_{2}}.$
@@
    Let $q_{1}\leq p_{1}<\infty$.
    1. a)
      Let $q_{2}<\infty$. Then
      $\leq\|a\|_{p_{1},q_{1}}^{q_{2}-q_{1}}\displaystyle\sum_{n=1}^{\infty}n^{-\frac{q_{2}-q_{1}}{p_{1}}}(a_{n}^{*})^{q_{1}}n^{\frac{q_{2}}{p_{2}}-1}$
      $=\|a\|_{p_{1},q_{1}}^{q_{2}-q_{1}}\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{2}}{p_{2}}-\frac{q_{2}}{p_{1}}}(a_{n}^{*})^{q_{1}}n^{\frac{q_{1}}{p_{1}}-1}$
      $\leq\|a\|_{p_{1},q_{1}}^{q_{2}-q_{1}}\displaystyle\sum_{n=1}^{\infty}(a_{n}^{*})^{q_{1}}n^{\frac{q_{1}}{p_{1}}-1}$
      $\leq\|a\|_{p_{1},q_{1}}^{q_{2}}.$
      Here the first inequality follows from the second part of Lemma 3.2 a_n^* ≤n^-1p1∥ a ∥_p_1,q_1, and we get the second inequality from the fact that n^q2p2 -q2p1¡1.
    2. b)
      Now let $q_{2}=\infty$ then:
      $\leq\displaystyle\sup_{n\in\mathbb{N}}n^{\frac{1}{p_{2}}}n^{-\frac{1}{p_{1}}}\|a\|_{p_{1},q_{1}}$
      $=\|a\|_{p_{1},q_{1}}\displaystyle\sup_{n\in\mathbb{N}}n^{\frac{1}{p_{2}}}n^{-\frac{1}{p_{1}}}$
      $\leq\|a\|_{p_{1},q_{1}}.$
      The first inequality follows from the second part of Lemma 3.2 a_n^* ≤n^-1p1∥ a ∥_p_1,q_1. And we get the second inequality from the fact that n^1p2 -1p1 ≤1.
@@
    Here the first inequality follows from the first part of Lemma 3.2 a_n^* ≤( q1p1)^1q1n^-1p1∥ a ∥_p_1,q_1 and we get the second inequality from the fact that n^q2p2 -q2p1 ≤1.
@@
    Let $p_{1}<q_{1}<\infty$ and $q_{1}\geq q_{2}$, then from the computations in the 3. part of the proof we have: ∥ a ∥_p_2,q_2≤( q1p1)^1q1 - 1q2∥ a ∥_p_1,q_1,
@@
Let $0<p_{1}\leq p_{2}\leq\infty$ and $q_{1}\in[1,\infty]$ and $q_{2}\in[1,\infty]$, then the norm of the embedding operator $I\colon\ell_{p_{1},q_{1}}\rightarrow\ell_{p_{2},q_{2}}$ is equal to:
@@
    Let $q_{1}\leq p_{1}$. We get the upper estimate of the operator norm for and $p_{1}<p_{2}$ from the first part of Lemma 3.4 and for and $p_{1}=p_{2}$ from the second part of Lemma 3.3. So $\|I\|\leq 1$. Now we will show that $I$ attains $1$ on the closed unit ball. For arbitrary $e^{j}$ we have that $(e^{j})^{*}=e^{1}$ and: ∥e^j∥_p_1,q_1 =(∑n=1 nq1p1- 1((enj)*)q1)^1q1 = (1^q1p1 - 11^q_1)^1q1 = 1. So $e^{j}\in B_{\ell_{p_{1},q_{1}}}$. Now we need to show that also $\|I(e^{j})\|_{p_{2},q_{2}}=\|e^{j}\|_{p_{2},q_{2}}=1$.
@@
    Let $p_{1}<q_{1}<\infty$ and $q_{1}\geq q_{2}$. We get the upper estimate for the operator norm from the fourth part of Lemma 3.4. So $\|I\|\leq 1$. We can show that $I$ attains $1$ on the closed unit ball at an arbitrary $e^{j}$. The computations are same as in the proof of 1. So again $\|I\|=1$.
@@
    Let $q_{1}=q_{2}=\infty$. We get the upper estimate for the operator norm from the fifth part of Lemma 3.4. So $\|I\|\leq 1$. We can show that $I$ attains $1$ on the closed unit ball at an arbitrary $e^{j}$. The computations are similar as in the proof of 1. ∥e^j∥_p_1,q_1 = supn∈N((enj)*)n1p1 = 1 = ∥ejp2,q2
@@
    Let $p_{1}<q_{1}=\infty$ and $q_{2}<\infty$. Once again we get the upper estimate for the operator norm from the second part of Lemma 3.4.
@@
    So indeed $I$ achieves it’s upper estimate on the closed unit ball and 4. is proven.
@@
Let $0<p<q<\infty$, then the norm of the embedding operator $I\colon\ell_{p,q}\rightarrow\ell_{p,\infty}$ is:
@@
We want to show that there exists a sequence $\{a^{n}\}_{n=1}^{\infty}$ of nonzero elements, where $a^{n}\in\ell_{p,q}$ for all $n\in\mathbb{N}$, such that:
@@
This would already imply that $\|I\|=\left(\frac{q}{p}\right)^{\frac{1}{q_{1}}}$, since we would be able to get arbitrarily close to our upper estimate $\left(\frac{q}{p}\right)^{\frac{1}{q}}$ of the operator norm $\|I\|$.
@@ 4. Examples
    We find the norm of operator $I$. Fix $a=\{a_{n}\}_{n=1}^{\infty}\in\ell_{p,q}$.
    1. a)
      Let $q=\infty$. Then $\|I(a)\|=\|a\|_{\infty}=\sup_{n\in\mathbb{N}}|a_{n}|\leq\sup_{n\in\mathbb{N}}\{n^{1/p}a_{n}^{*}\}=\|a\|_{p,q}$, where the inequality results from the fact that $n^{1/p}\geq 1$ for all $n\in\mathbb{N}$ and the Definition 1.6 of the decreasing rearrangement of $a$.
    2. b)
      Let $q<\infty$. Then $\|I(a)\|=\|a\|_{\infty}=\sup_{n\in\mathbb{N}}|a_{n}|\leq\left(\displaystyle\sum_{n=1}^{\infty}(a_{n}^{*})^{q}n^{\frac{q}{p}-1}\right)^{\frac{1}{q}}=\|a\|_{p,q}$, where the inequality results from the Definition 1.6 of the decreasing rearrangement of $a$ and because $\sup_{n\in\mathbb{N}}|a_{n}|=a_{1}^{*}=1^{\frac{q}{p}-1}(a_{1}^{*})^{q}$.
    Thus, $\|I\|\leq 1$ for any $p\in[1,\infty)$, $q\in[1,\infty]$. Now we will show that $I$ achieves $1$ on the closed unit ball. For an arbitrary $e^{j}$ we have $(e^{j})^{*}=e^{1}$, therefore $\|e^{j}\|_{p,q}=\|e^{1}\|_{p,q}=1$, thus $e^{j}\in B_{\ell_{p,q}}$ and also $\|I(e^{j})\|_{\infty}=\|e^{j}\|_{\infty}=1$, hence $\|I|\|=1$.
So, from Theorem 2.3 the embedding operator $I$ is maximally non-compact and $\alpha(I)=\|I\|=1$. ∎
@@
Let $0<p_{1}<p_{2}\leq\infty$, $q_{1},q_{2}\in[1,\infty]$, $\min\{p_{2},q_{2}\}<\infty$ and $\ell^{p_{1},q_{1}}$, $\ell^{p_{2},q_{2}}$ Lorentz sequence spaces. Then the embedding operator $I\colon\ell^{p_{1},q_{1}}\rightarrow\ell^{p_{2},q_{2}}$ is maximally non-compact.
@@
So, the prerequisites of Theorem 2.3 are satisfied, the embedding operator $I$ is maximally non-compact and $\alpha(I)=\|I\|$, where the values of $\|I\|$ for some individual relations between $p_{1}$, $q_{1}$ and $q_{2}$ are computed in Theorem 3.5 or Theorem 3.6. ∎
@@
We have $\ell^{p,q}\subset c_{0}\subset\ell^{\infty}$ from Lemma 3.1 and I=1||I||=1 follows from the computations in Example 4.1
@@
    Now when sending $\varepsilon\rightarrow 0_{+}$ we get our claim. Therefore σ_ℓ_p,q ≤σ¡ 2∥I∥ and all the prerequisites of Theorem 2.4 are satisfied, so $I$ is not maximally non-compact and $\alpha(I)\leq\sigma/2<1=\|I\|$.
@@
    where the second inequality follows from our prerequisite $p\leq q$ thus $n^{\frac{q}{p}-1}\geq 1$. and the fourth follows from the Lemma 1.15
    Now when sending $\varepsilon\rightarrow 0_{+}$ we get our claim. Therefore σ_ℓ_p,q ≤1 ¡ 2∥I∥ and all the prerequisites of Theorem 2.4 are satisfied, so $I$ is not maximally non-compact and $\alpha(I)\leq 1/2<1=\|I\|$.
@@
    We computed that $\sigma_{\ell_{p,q}}\leq\sigma=1+2^{-1/p}<2\|I\|$. Now again from Theorem 2.4 we get that $I$ is not maximally non-compact and $\alpha(I)\leq\varrho=2^{-1}(1+2^{-1/p})<1$.
@@
    Now let $q<p$. By the second part of Theorem 3.3 $\ell_{p,q}\hookrightarrow\ell_{p,\infty}$ and $\|y\|_{p,\infty}\leq\|y\|_{p,q}$ for all $y\in\ell_{p,q}$. Now assume $y\in B_{\ell^{p,q}}$. Then $y\in B_{\ell^{p,\infty}}$, so from the proof of the previous case, where $q=\infty$, we again get $y_{n}^{*}\leq n^{-1/p}$ and $\sigma_{\ell_{p,q}}\leq 1+2^{-1/p}<2\|I\|$. Now again from Theorem 2.4 we get that $I$ is not maximally non-compact and α(I) ≤ϱ= 2^-1(1+2^-1/p) ¡ 1.

### v2

@@ 3. Lorentz sequence spaces
Let us assume that $\ell^{p,q}\not\subset c_{0}$ and take $a\in\ell^{p,q}\setminus c_{0}$. Then there exists $\varepsilon>0$, such that for all $n_{0}\in\mathbb{N}$ there exists $n\geq n_{0},\ n\in\mathbb{N}\colon|a_{n}|\geq\varepsilon$, which implies that $a_{n}^{*}\ \geq\ \varepsilon$ for every $n\in\mathbb{N}$. First, let us look at the case when $q<\infty$ and $p\in(0,\infty]$. Then:
@@
The statements of the following two lemmas occur in the article [10, p. 77-78], however, the verification of the statements is only sketched there. Some ideas of proofs are furthermore similar to analogous arguments known from the nonatomic case (see e.g. [15, Section 8.2]), but since in the discrete case other techniques have to be used, we include the simple verification.
@@
Let $p\in(0,\infty]$, $q\in(0,\infty)$ and $a=\{a_{n}\}_{n=1}^{\infty}\in\ell^{p,q}$, then for all $n\in\mathbb{N}$ it holds that
@@
    Here, the second inequality follows from the fact that $(a_{i}^{*})$ is the decreasing rearrangement. The third inequality results from the Lagrange mean value theorem. Indeed, for all $i\in\mathbb{N}$, there exists $\xi\in(i-1,i)$, that satisfies for the function $f(t)=\frac{p}{q}t^{\frac{q}{p}}$, where $t\geq 0$ following equality:
    $f^{{}^{\prime}}(\xi)=\xi^{\frac{q}{p}-1}=\frac{p}{q}\left[i^{\frac{q}{p}}-(i-1)^{\frac{q}{p}}\right].$
    Also $i^{\frac{q}{p}-1}\geq\xi^{\frac{q}{p}-1}$, because $\frac{q}{p}\geq 1$ and $i>\xi>0$. From the estimates above we get the third inequality and 1. is proven.
@@
    Here the second inequality follows from the fact that $\frac{q}{p}\in[0,1)$ and $i\leq n$. Therefore $\left(\frac{q}{p}-1\right)\in[-1,0)$, so we get $n^{\left(\frac{q}{p}-1\right)}\leq i^{\left(\frac{q}{p}-1\right)}$, and 2. is proven.
@@
$\ell^{p,q_{1}}\hookrightarrow\ell^{p,q_{2}},$
and for all $a\in\ell^{p,q_{1}}$ it holds that
@@
    $\|a\|_{p,q_{2}}\leq\left(\frac{q_{1}}{p}\right)^{\frac{1}{q_{1}}-\frac{1}{q_{2}}}\|a\|_{p,q_{1}}$ if $p<q_{1}$,
@@
    Let $p<q_{1}$ and $q_{2}\in(q_{1},\infty]$. We compute:
    $\|a\|_{p,q_{2}}$ $=\left\|n^{\frac{1}{p}-\frac{1}{q_{2}}}a_{n}^{*}\right\|_{q_{2}}$
    $=\left\|n^{\frac{1}{p}-\frac{1}{q_{2}}}(a_{n}^{*})^{1-\frac{q_{1}}{q_{2}}}(a_{n}^{*})^{\frac{q_{1}}{q_{2}}}\right\|_{q_{2}}$
    $\leq\left\|\left(\frac{q_{1}}{p}\right)^{\frac{1}{q_{1}}-\frac{1}{q_{2}}}\|a\|_{p,q_{1}}^{1-\frac{q_{1}}{q_{2}}}n^{-\frac{1}{p}+\frac{q_{1}}{pq_{2}}}n^{\frac{1}{p}-\frac{1}{q_{2}}}(a_{n}^{*})^{\frac{q_{1}}{q_{2}}}\right\|_{q_{2}}$
    $=\left(\frac{q_{1}}{p}\right)^{\frac{1}{q_{1}}-\frac{1}{q_{2}}}\|a\|_{p,q_{1}}^{1-\frac{q_{1}}{q_{2}}}\left\|n^{\frac{q_{1}}{pq_{2}}-\frac{1}{q_{2}}}(a_{n}^{*})^{\frac{q_{1}}{q_{2}}}\right\|_{q_{2}}$
    $=\left(\frac{q_{1}}{p}\right)^{\frac{1}{q_{1}}-\frac{1}{q_{2}}}\|a\|_{p,q_{1}}^{1-\frac{q_{1}}{q_{2}}}\left\|n^{\frac{1}{p}-\frac{1}{q_{1}}}a_{n}^{*}\right\|_{q_{1}}^{\frac{q_{1}}{q_{2}}}$
    $=\left(\frac{q_{1}}{p}\right)^{\frac{1}{q_{1}}-\frac{1}{q_{2}}}\|a\|_{p,q_{1}}.$
    Here the first inequality again follows from the first part of Lemma 3.2. For the case $q_{2}=\infty$ we apply the convention $1/q_{2}=1/\infty=0$.
@@
    Now, let $p\geq q_{1}$ and $q_{2}\in(q_{1},\infty]$. Then, the following inequality follows from the second part of Lemma 3.2 and the computations are analogous to the first case:
    $\|a\|_{p,q_{2}}$ $=\left\|n^{\frac{1}{p}-\frac{1}{q_{2}}}(a_{n}^{*})^{1-\frac{q_{1}}{q_{2}}}(a_{n}^{*})^{\frac{q_{1}}{q_{2}}}\right\|_{q_{2}}$
    $\leq\|a\|_{p,q_{1}}^{1-\frac{q_{1}}{q_{2}}}\left\|n^{\frac{q_{1}}{pq_{2}}-\frac{1}{q_{2}}}(a_{n}^{*})^{\frac{q_{1}}{q_{2}}}\right\|_{q_{2}}$
    $=\|a\|_{p,q_{1}}.$
@@
    Let $q_{1}\leq p_{1}<\infty$ and $q_{2}\in(0,\infty]$, then
    $\|a\|_{p_{2},q_{2}}$ $=\left\|n^{\frac{1}{p_{2}}-\frac{1}{q_{2}}}(a_{n}^{*})^{1-\frac{q_{1}}{q_{2}}}(a_{n}^{*})^{\frac{q_{1}}{q_{2}}}\right\|_{q_{2}}$
    $\leq\|a\|_{p_{1},q_{1}}^{1-\frac{q_{1}}{q_{2}}}\left\|n^{\frac{1}{p_{2}}-\frac{1}{p_{1}}}n^{\frac{q_{1}}{p_{1}q_{2}}-\frac{1}{q_{2}}}(a_{n}^{*})^{\frac{q_{1}}{q_{2}}}\right\|_{q_{2}}$
    $\leq\|a\|_{p_{1},q_{1}}^{1-\frac{q_{1}}{q_{2}}}\left\|n^{\frac{q_{1}}{p_{1}q_{2}}-\frac{1}{q_{2}}}(a_{n}^{*})^{\frac{q_{1}}{q_{2}}}\right\|_{q_{2}}$
    $=\|a\|_{p_{1},q_{1}}^{1-\frac{q_{1}}{q_{2}}}\left\|n^{\frac{1}{p_{1}}-\frac{1}{q_{1}}}a_{n}^{*}\right\|_{q_{1}}^{\frac{q_{1}}{q_{2}}}$
    $=\|a\|_{p,q_{1}}.$
    Here the first inequality follows from the second part of Lemma 3.2
    $a_{n}^{*}\leq n^{-\frac{1}{p_{1}}}\|a\|_{p_{1},q_{1}},$
    and we get the second inequality from the fact that
    $n^{\frac{1}{p_{2}}-\frac{1}{p_{1}}}\leq 1.$
@@
    Here the first inequality follows from the first part of Lemma 3.2
    $a_{n}^{*}\leq\left(\frac{q_{1}}{p_{1}}\right)^{\frac{1}{q_{1}}}n^{-\frac{1}{p_{1}}}\|a\|_{p_{1},q_{1}}$
    and we get the second inequality from the fact that
    $n^{\frac{q_{2}}{p_{2}}-\frac{q_{2}}{p_{1}}}\leq 1.$
@@
    Let $p_{1}<q_{1}<\infty$ and $q_{1}\geq q_{2}$, then from the computations in the 3. part of the proof we have:
    $\|a\|_{p_{2},q_{2}}\leq\left(\frac{q_{1}}{p_{1}}\right)^{\frac{1}{q_{1}}-\frac{1}{q_{2}}}\|a\|_{p_{1},q_{1}},$
@@
Let $0<p_{1}\leq p_{2}\leq\infty$ and $q_{1},q_{2}\in(0,\infty]$. If $p_{1}=p_{2}$, let us also assume $q_{1}<q_{2}$. Then the norm of the embedding operator $I\colon\ell^{p_{1},q_{1}}\rightarrow\ell^{p_{2},q_{2}}$ is equal to:
@@
    Let $q_{1}\leq p_{1}$. We get the upper estimate of the operator norm for and $p_{1}<p_{2}$ from the first part of Theorem 3.4 and for and $p_{1}=p_{2}$ from the second part of Lemma 3.3. So $\|I\|\leq 1$. Now we will show that $I$ attains $1$ on the closed unit ball. For arbitrary $e^{j}$ we have that $(e^{j})^{*}=e^{1}$ and:
    $\|e^{j}\|_{p_{1},q_{1}}=\left(\displaystyle\sum_{n=1}^{\infty}n^{\frac{q_{1}}{p_{1}}-1}((e_{n}^{j})^{*})^{q_{1}}\right)^{\frac{1}{q_{1}}}=\left(1^{\frac{q_{1}}{p_{1}}-1}1^{q_{1}}\right)^{\frac{1}{q_{1}}}=1.$
    So $e^{j}\in B_{\ell^{p_{1},q_{1}}}$. Now we need to show that also $\|I(e^{j})\|_{p_{2},q_{2}}=\|e^{j}\|_{p_{2},q_{2}}=1$.
@@
    Let $p_{1}<q_{1}<\infty$ and $q_{1}\geq q_{2}$. We get the upper estimate for the operator norm from the fourth part of Theorem 3.4. So $\|I\|\leq 1$. We can show that $I$ attains $1$ on the closed unit ball at an arbitrary $e^{j}$. The computations are same as in the proof of 1. So again $\|I\|=1$.
@@
    Let $q_{1}=q_{2}=\infty$. We get the upper estimate for the operator norm from the fifth part of Theorem 3.4. So $\|I\|\leq 1$. We can show that $I$ attains $1$ on the closed unit ball at an arbitrary $e^{j}$. The computations are similar as in the proof of 1.
    $\|e^{j}\|_{p_{1},q_{1}}=\displaystyle\sup_{n\in\mathbb{N}}((e_{n}^{j})^{*})n^{\frac{1}{p_{1}}}=1=\|e^{j}\|_{p_{2},q_{2}}$
@@
    Let $p_{1}<q_{1}=\infty$ and $q_{2}<\infty$. Once again we get the upper estimate for the operator norm from the second part of Theorem 3.4.
@@
    $a=a^{*}=(1,2^{-\frac{1}{p_{1}}},3^{-\frac{1}{p_{1}}},\dots,n^{-\frac{1}{p_{1}}},\dots)$, then
    $\|a\|_{p_{1},\infty}=\displaystyle\sup_{n\in\mathbb{N}}a_{n}^{*}n^{\frac{1}{p_{1}}}=\displaystyle\sup_{n\in\mathbb{N}}n^{-\frac{1}{p_{1}}}n^{\frac{1}{p_{1}}}=1.$
@@
    So indeed $I$ achieves its upper estimate on the closed unit ball and 4. is proven.
In the case 3 of Theorem 3.4, we do not know whether the constant is optimal in full generality, except in the case treated in the next result.
@@
Let $0<p<q<\infty$, then the norm of the embedding operator $I\colon\ell^{p,q}\rightarrow\ell^{p,\infty}$ is:
@@
We want to show that there exists a sequence $\{a^{n}\}_{n=1}^{\infty}$ of nonzero elements, where $a^{n}\in\ell^{p,q}$ for all $n\in\mathbb{N}$, such that:
@@
This would already imply that $\|I\|=\left(\frac{q}{p}\right)^{\frac{1}{q}}$, since we would be able to get arbitrarily close to our upper estimate $\left(\frac{q}{p}\right)^{\frac{1}{q}}$ of the operator norm $\|I\|$.
@@ 4. Examples
    Since $\|\cdot\|_{c_{0}}=\|\cdot\|_{\infty}=\|\cdot\|_{\infty,\infty}$, we already know that $\|I\|=1$ from Theorem 3.5. We now have $\ell^{p,q}\hookrightarrow c_{0}$.
So, from Theorem 2.3 the embedding operator $I$ is maximally non-compact and $\alpha(I)=1$. ∎
@@
Let $0<p\leq\infty$, $0<q_{1}<q_{2}\leq\infty$, $\min\{p,q_{2}\}<\infty$, and let $\ell^{p_{1},q_{1}}$, $\ell^{p_{2},q_{2}}$ be Lorentz sequence spaces. Then the embedding operator $I\colon\ell^{p,q_{1}}\rightarrow\ell^{p,q_{2}}$ is maximally non-compact.
Proof.
We have
  1. 1.
    $\ell^{p,q_{1}}\hookrightarrow\ell^{p,q_{2}}$ from Theorem 3.3,
  2. 2.
    $\ell^{p,q_{1}}$, $\ell^{p,q_{2}}\subset c_{0}$ from Lemma 3.1.
So, the prerequisites of Theorem 2.3 are satisfied, the embedding operator $I$ is maximally non-compact and $\alpha(I)=\|I\|$, where the values of $\|I\|$ for some individual relations between $p$, $q_{1}$ and $q_{2}$ are computed in Theorem 3.5 or Theorem 3.6. ∎
Proposition 4.3.
Let $0<p_{1}<p_{2}\leq\infty$, $q_{1},q_{2}\in(0,\infty]$, $\min\{p_{2},q_{2}\}<\infty$, and let $\ell^{p_{1},q_{1}}$, $\ell^{p_{2},q_{2}}$ be Lorentz sequence spaces. Then the embedding operator $I\colon\ell^{p_{1},q_{1}}\rightarrow\ell^{p_{2},q_{2}}$ is maximally non-compact.
@@
So, the prerequisites of Theorem 2.3 are satisfied, the embedding operator $I$ is maximally non-compact and $\alpha(I)=\|I\|$, where the values of $\|I\|$ for some individual relations between $p_{1}$, $q_{1}$ and $q_{2}$ are computed in Theorem 3.5. ∎
@@
We have $\ell^{p,q}\subset c_{0}\subset\ell^{\infty}$ from Lemma 3.1 and I=1||I||=1 follows from the fact that $\|\cdot\|_{\infty}=\|\cdot\|_{\infty,\infty}$ and Theorem 3.5
@@
    Now when sending $\varepsilon\rightarrow 0_{+}$ we get our claim. Therefore
    $\sigma_{\ell^{p,q}}\leq\sigma<2\|I\|$
    and all the prerequisites of Theorem 2.4 are satisfied, so $I$ is not maximally non-compact and $\alpha(I)\leq\sigma/2<1=\|I\|$.
@@
    Now when sending $\varepsilon\rightarrow 0_{+}$ we get our claim. Therefore
    $\sigma_{\ell^{p,q}}\leq 1<2\|I\|$
    and all the prerequisites of Theorem 2.4 are satisfied, so $I$ is not maximally non-compact and $\alpha(I)\leq 1/2<1=\|I\|$.
@@
    We computed that $\sigma_{\ell^{p,q}}\leq\sigma=1+2^{-1/p}<2\|I\|$. Now again from Theorem 2.4 we get that $I$ is not maximally non-compact and $\alpha(I)\leq\varrho=2^{-1}(1+2^{-1/p})<1$.
@@
    Now let $q<p$. By the second part of Theorem 3.3 $\ell^{p,q}\hookrightarrow\ell_{p,\infty}$ and $\|y\|_{p,\infty}\leq\|y\|_{p,q}$ for all $y\in\ell^{p,q}$. Now assume $y\in B_{\ell^{p,q}}$. Then $y\in B_{\ell^{p,\infty}}$, so from the proof of the previous case, where $q=\infty$, we again get $y_{n}^{*}\leq n^{-1/p}$ and $\sigma_{\ell^{p,q}}\leq 1+2^{-1/p}<2\|I\|$. Now again from Theorem 2.4 we get that $I$ is not maximally non-compact and
    $\alpha(I)\leq\varrho=2^{-1}(1+2^{-1/p})<1.$
@@
The author would also like to thank both anonymous referees for their thorough and critical reading of the paper and for plenty of very valuable remarks and suggestions for adjustments and further revision.